
Normal distribution
Presentation
•
Mathematics
•
11th Grade
•
Practice Problem
•
Medium
Standards-aligned
Rizki Alfath
Used 14+ times
FREE Resource
17 Slides • 15 Questions
1
2
3
4
5
6
7
Multiple Choice
Given that:
μ = 56 and σ = 7
Find the standardized score for a raw score of 70.
z = 0
z = 1.5
z = –0.5
z = 2
8
Multiple Choice
Given that:
μ = 56 and σ = 7
Find the standardized score for a raw score of 52.5.
z = 0
z = 1.5
z = –0.5
z = 2
9
Multiple Choice
Given that:
μ = 56 and σ = 7
Find the standardized score for a raw score of 66.5.
z = 0
z = 1.5
z = –0.5
z = 2
10
Multiple Choice
Given that:
μ = 56 and σ = 7
Find the standardized score for a raw score of 56.
z = 0
z = 1.5
z = –0.5
z = 2
11
Multiple Choice
Given that:
μ = 56 and σ = 7
Find the a raw score for standardized score of 1.3.
x = 65.1
x = 39.2
x = 53.2
x = 70
12
Multiple Choice
Given that:
μ = 56 and σ = 7
Find the a raw score for standardized score of –2.4.
x = 65.1
x = 39.2
x = 53.2
x = 70
13
Multiple Choice
Given that:
μ = 56 and σ = 7
Find the a raw score for standardized score of –0.4.
x = 65.1
x = 39.2
x = 53.2
x = 70
14
Multiple Choice
Given that:
μ = 56 and σ = 7
Find the a raw score for standardized score of 2.0.
x = 65.1
x = 39.2
x = 53.2
x = 70
15
16
17
18
19
20
If Z ~ N(0, 12) find P(Z < 0.247).
P(Z < 0.247) = 0.5948 + 0.0027 = 0.5975
21
Multiple Choice
If Z ~ N(0, 12) find Φ(1.62).
[Hint: P(Z < 1.62) = Φ(1.62)]
0.9474
0.2236
0.0934
0.8438
22
Multiple Choice
If Z ~ N(0, 12) find P(Z > 0.76).
[Hint: P(Z > 0.76) = 1 – Φ(0.76)]
0.9474
0.2236
0.0934
0.8438
23
Multiple Choice
If Z ~ N(0, 12) find P(Z < –1.32).
[Hint: P(Z < –1.32) = 1 – Φ(1.32)]
0.9474
0.2236
0.0934
0.8438
24
Multiple Choice
If Z ~ N(0, 12) find P(–1.2 < Z < 1.738).
[Hint: P(–1.2 < Z < 1.738) = Φ(1.738) – Φ(–1.2)]
0.9474
0.2236
0.0934
0.8438
25
If X ~ N(2, 52) find P(X < 7).
26
Fill in the Blanks
If X ~ N(2, 52) find P(X < 9).
27
If X ~ N(2, 52) find P(X > 11).
28
Fill in the Blanks
If X ~ N(2, 52) find P(X > 10).
29
If X ~ N(2, 52) find P(|X| < 3).
30
Fill in the Blanks
If X ~ N(2, 52) find P(X > 10).
31
If X ~ N(2, 52) find P(|X - 2| < 6).
32
If X ~ N(2, 52).
Find x such that P(X > x) = 0.05
Show answer
Auto Play
Slide 1 / 32
SLIDE
Similar Resources on Wayground
20 questions
Persamaan Lingkaran
Presentation
•
11th Grade
20 questions
Graphing Polynomials
Presentation
•
11th Grade
23 questions
Unit 3: Mid-Chapter Review
Presentation
•
10th - 12th Grade
27 questions
Exponential Function Lesson
Presentation
•
11th Grade
24 questions
4.26_AMDM_Simple & Compound Interest
Presentation
•
10th - 12th Grade
21 questions
Composition of Functions
Presentation
•
11th Grade
25 questions
Transformasi geometri (Translasi dan Refleksi)
Presentation
•
11th Grade
20 questions
Reading Graphs
Presentation
•
11th Grade
Popular Resources on Wayground
10 questions
Safety Day 2026 – Student Review
Quiz
•
9th Grade
21 questions
Continents and Oceans
Quiz
•
6th Grade
20 questions
Parts of Speech
Quiz
•
5th Grade
16 questions
Subject & Predicate
Quiz
•
5th Grade
25 questions
Multiplication Facts
Quiz
•
5th Grade
12 questions
Map Skills
Quiz
•
3rd Grade
22 questions
Continents and Oceans
Quiz
•
5th Grade
15 questions
Rounding Decimals
Quiz
•
5th Grade
Discover more resources for Mathematics
12 questions
Identify Rigid Transformations
Quiz
•
8th - 12th Grade
10 questions
Complementary, Supplementary, Vertical, and Linear Pairs
Quiz
•
9th - 12th Grade
22 questions
Identify Vertical and Horizontal Translations
Quiz
•
9th - 12th Grade
10 questions
Graphing Inequalities on a Number Line
Quiz
•
9th - 12th Grade
17 questions
Identifying Parallel And Perpendicular Lines
Quiz
•
9th - 12th Grade
18 questions
Integer Operations
Quiz
•
5th - 12th Grade
18 questions
Parallel lines cut by a transversal
Quiz
•
9th - 11th Grade
14 questions
Evaluate Linear Functions and Their Applications
Quiz
•
9th - 12th Grade